OpenStudy (anonymous):

What is the half life of a first-order reaction with a rate constant of 2.43*10^-2 min^-1 ?

OpenStudy (anonymous):

The decay constant, λ, equals ln(2)/half life. So; \[\lambda = \frac {\ln(2)}{t_{1/2}} \] \[t_{\frac{1}{2}} = \frac{\ln(2)}{2.43*10^{-2}} = 28.52 \min\]